Determine whether or not the reaction indicated will occur spontaneously. 2Cu+(aq) + Ni(s) Ú 2Cu(s) + Ni2+(aq)
Answer #1

The Given Reaction:

2Cut (aq) + Ni(s) + 2Cu(s) + Ni2+ (aq)

Half cell Reaction:

Cathode

Cut(aq) + le + Cu(s)   ECut/Cu = 0.53V....1

Anode

Ni(s) + Ni2+ + 2eEN 2+/Ni = -0.25V........2

Ecell = ECu2+/Cu – Epi2+/Ni= 0.53 -(-0.25) = 0.78= 0.53 - (-0.25) = +0.78V

Here, E_{cell}^{o} valve is +ve, hence the cell will be feasible and Ni will displace copper from its salt solution. In other words Nickel will reduce Copper.

And AG = -FE

If E^{o} valve is +ve the the valve of \Delta G^{o} will be negative, we can say that the reaction will occur spontaneously.